OpenStudy (anonymous):

Janet's gross pay is $750, from which $31 is deducted for OASDI, $10 for Medicare, and $14 for income tax. What is her net pay?

OpenStudy (anonymous):

A. $695 B. $709 C. $719

ganeshie8 (ganeshie8):

subtract deductions from the gross pay

ganeshie8 (ganeshie8):

deductions = 31 + 10 + 14 = ?

OpenStudy (anonymous):

subtract deductions from the gross pay like (deductions = 31 + 10 + 14) =??????????????????????????